the speed of a train a goes 10mph slower then the speed of train b train a traveles 210 miles and train b traveles 260 miles find the speed of the trains?
Answer by mananth(16949)

the speed of a train a goes 10mph slower then the speed of train b train a traveles 210 miles and train b traveles 260 miles find the speed of the trains?
...
train B x mph
train A x-10 mph
..
210/(x-10)= 260/x
210x=260(x-10)
210x=260x-2600
50x = 2600
/50
x= 2600/50
x=52mph train B
train A = 52-10=42 mph
...
m.ananth@hotmail.ca